NAOJ GW Elog Logbook 3.2

Since the oplev beam reflected by the input mirror was very high as reported in elog2798, I tweaked a steering mirror just after the oplev laser source for input mirror (Fig. 1). The oplev sum for input gets around 5000.
Fig. 2 shows the oplev spectrum of input mirror (top right). There are large peaks at 2 Hz and 3.4 Hz. We need to open the input chamber and check the suspension.

[Aritomi, Yuhang, Michael]
We found that the green reflection from input mirror is very high and hits on the upper edge of gate valve between BS/input chambers. The oplev light for input mirror is also very high as shown in the attached picture. We tried to move input mirror with picomotor, but the input mirror did not move. We will align oplev and measure the input oplev spectrum tomorrow.

I connected the END YAW to the motor A of driver although I don't know the motor A is working or not.

I took a function generator and an oscilloscope from filter cavity clean room to modulate EOM and acquire data in ATC.
The RF PD is moved to OPO reflection to acquire PDH signal. A minicircuit zx05-1l-s+ is used for RF signal demodulation. 10dBm 40MHz signal is sent to New focus 4003 EOM to get a phase modulation around 0.07~0.21rad. 3dBm signal is used as LO. We got error signal as attached figure. Some glitches are found and we will try to understand it better tomorrow.
One possible reason is that I should put a DC block between RF PD and mixer. Note that demodulation phase is not optimized yet.

To measure effectively the internal optical losses of OPO, we need to inject laser from the in-coupling mirror side (elog2784).
Therefore, I flipped OPO and moved OPO closer to 75mm lens as described in elog2790. However, I found that moving 54mm is far too much, whose reason is not understood yet. After moving OPO farther from 75mm lens, the mode matching situation is better. However, I found that the beam is going into the in-coupling mirror but close to the edge of its hole. So I loosed in-coupling mirror and moved it vertically down. After this, I found it diffcult to recover alignment at the first glance.
However, I realized that it is actually very easy to re-align OPO using transmission camera and injection/reflection beam overlaping (without using periscope/translation stage). The method is bascially the same with the procedure provided in elog2783 apart from different beam parameters and almost not seperating the alignment of crystal and in-coupling mirror but aligning them as almost a whole. And I found it is possible to choose a personal preferred position of in-coupling mirror and just fix it with the crystal side of OPO. Then align them as a whole.
After this easier alignment procudure, checking by eye, I found the laser is far enough from the edge of OPO two sides holes. Then I aligned OPO until mode matching level is around 95%. Now OPO is ready for internal optical losses measurement.

[Aritomi, Yuhang]
After the evacuation of PR/BS chambers, we aligned PR/BS picomotors to center the PR reference and the first target. Then we checked the second target. Although we scanned BS alignment a lot, we could see only scattered light on the second target.
Instead of the nominal PR reference, we used the green beam position in the gate valve between BS/input chambers as PR reference (Fig.1). We centered the green beam in the gate valve between BS/input chambers by PR picomotor and centered the green beam at first target by BS picomotor. Then we found the green beam in the second target. Since the green beam position in the nominal PR reference changed, we made a new PR reference as Fig.2.

We firstly used Rotary Vane Vacuum Pump (Alcatel 2100A) to evacuate. The speed is controlled to be less than 1mbar/s to avoid large pressure applied on in-vacuum components. Its ultimate achievable pressure is 0.03mbar.
Since the turbo pump (STP1003) requires outlet pressure smaller than 0.1torr (0.13mbar), we waited for rotary pump to evacuate PR/BS chambers until pressure reached 0.1mbar.
The small rotary pump and turbo pump were kept on with gate valve closed between turbo pump and PR/BS chambers. So we just opened the gate valve after PR/BS chambers pressure is lower than 0.1mbar.
We will keep turbo pump on until the PR/BS chambers pressure reaches 1e-6 mbar. Then we will open the gate valve between PR/BS chambers and input mirror chamber.

[Takahashi, Aritomi, Yuhang]
Since the BS picomotor did not move, we opened the BS chamber and checked the BS picomotor. The BS picomotors for pitch and yaw were connected to the motors A and B in driver 1, respectively. However, we found that the motors A and B in driver 1 were broken. We decided to use the motor C in driver 1 for both BS pitch and yaw temporarily.
We noticed that the BS picomotor did not move in yaw direction although we heard the sound of the picomotor. Takahashi-san found that it was because the BS was touching the stopper of the mirror. After Takahashi-san fixed it, the BS picomotor could move in yaw.
After we adjusted the BS picomotor, the green beam is hitting on center of the first target. Finally, we started to evacuate the PR/BS chambers.
We firstly used Rotary Vane Vacuum Pump (Alcatel 2100A) to evacuate. The speed is controlled to be less than 1mbar/s to avoid large pressure applied on in-vacuum components. Its ultimate achievable pressure is 0.03mbar.
Since the turbo pump (STP1003) requires outlet pressure smaller than 0.1torr (0.13mbar), we waited for rotary pump to evacuate PR/BS chambers until pressure reached 0.1mbar.
The small rotary pump and turbo pump were kept on with gate valve closed between turbo pump and PR/BS chambers. So we just opened the gate valve after PR/BS chambers pressure is lower than 0.1mbar.
We will keep turbo pump on until the PR/BS chambers pressure reaches 1e-6 mbar. Then we will open the gate valve between PR/BS chambers and input mirror chamber.

Michael and Yuhang
This Friday (20220114), we changed optical set-up from what described in elog2788. This is because we found one of the steering mirrors in front of OPO cannot effectively align laser beam inside OPO. We checked beam parameters in elog2486, which shows that the non-effective steering mirror is located just around the beam waist after 40mm lens. This explains why this mirror shows anomaly while OPO alignment procedure.
To solve this problem, we modified the optical set-up. We add one steering mirror just before OPO while leaving enough space for a BS between it and OPO (Fig.1). To confirm this set-up is capable of aligning beam and the OPO still has good internal alignment, we aligned OPO while laser beam is injected from the crystal side of OPO. We found the newly add steering mirror helps to optimize alignment.
In the end, we achieved a decent alignment by moving the position of lenses. The TEM02 mode is optimized by moving position of lenses. The position of lens are shown in Fig.2 and Fig.3.
We also characterize the mode matching level by measuring the TEM00, TEM01 and TEM02 modes. Their seperate power is shown in Fig.4, 5, and 6. This indicates a mode-mismatch of (5.6+18.8)/(5.6+18.8+282) = 8%
Using parameters of OPO, we simulate the beam parameters as Fig.7. From this simulation, we see the beam waist should be located in front of crystal when injected from the crystal side. But if we inject from in-coupling mirror side, the beam waist should be located inside OPO. Therefore, to measure optical losses and flip OPO housing, we need to flip OPO and move it OPO by 54mm closer to 75mm lens.

[Aritomi, Yuhang]
Last year, we found that the BS picomotor in yaw direction does not move as reported in elog2767.
Today, when we tried to move the BS picomotor in pitch direction with step of 50, the picomotor moved a lot and got stuck. We will open BS chamber to check the picomotor next Monday.

Michael and Yuhang
As discussed in elog2784, we need to inject laser from OPO in-coupling mirror side to distinguish different OPO internal losses values. This is the reason of work reported in this elog. To achieve the measurement, we have done the followings:
1. Remove periscope and rotation stage for OPO
2. Prepare a 50:50 BS placing in front of OPO to extract OPO reflection signal information
3. Lower the height of OPO transmission BS, PD and camera
(these works are shown in Fig.1)
4. Confirm and adjust the flatness of laser beam after removing periscope
(this work is shown in Fig.2,3,4)
5. Put OPO in place and optimize its position by hand. Try to get a reasonable OPO transmission mode. We have found a TEM01 mode now.
Next step: Find OPO transmission signal on PD. Optimize OPO injection beam alignment.

The camera used in this set-up is found to be tilted today. Fig.1 shows that this tilt is introduced through the interface between camera and post.
During the alignment, we use camera to characterize cavity HG modes shapes. When the beam and camera height is aligned, we found beam appears not in the center of screen as Fig.2. We suspect the tilt between camera and post is the cause of the cavity modes mis-center as Fig.2.
Note that Fig.2 shows the non-ideal alignment of crystal, which has a black defect.

This entry reports the measurement of polarization angle and birefringence of the shinkosha 7 sample measured in the same orientation as Manuel absorption measurement (ie arrow at the top pointing towards the imaging unit).
It is in quite good agreement with the previous measurement.
Note that the birefringence results reported here neglect the angle of incidence of the pump beam

Considering the CCFC+green OLTF shown in elog2758 and the IR locking accuracy (CCFC error signal) without CCFC, we can estimate the expected IR locking accuracy with CCFC. I compared the estimated IR locking accuracy with CCFC and the measured one as shown in the attached figure. As you can see, the two spectra agree well. The discrepancy between 10-100 Hz could be because the measured spectrum would be limited by the spectrum analyzer noise between 10-100 Hz.

Since OPO is finally closed, the next step is to characterize the intra-cavity losses. This is important for us because we are suspecting some of the optical losses are from OPO (current estimated loss budget for FDS). So this is an important step to understand the loss budget in the frequency dependent squeezing experiment.
I modified a bit the code to see the difference of measurement for different OPO intra-cavity losses.
Now, the laser is injected from the crystal side of OPO. I did a simulation of this case as Fig. 1. In this case, we will miss the information of OPO reflection. The blue and orange curves overlap for reflection.
If laser is injected from in-coupling mirror, as shown in Fig.2, we find that although decay time is not enough to indicate optical losses. We can extract losses from reflection signal.
So we will rotate OPO next week and inject laser from the in-coupling mirror side.

Michael and Yuhang
After the in-coupling mirror cleaning as reported in elog2774 and elog2776, we tried to close OPO again. Apart from mirror cleaning, this time, we tried to align OPO better than elog2769. To make sure a good crystal alignment, we paid attention to:
1. Making sure the flatness of laser beam in pitch/yaw direction using alignment tools such as rulers. Use power of few milliwat now.
2. Change to less than microwatt. Making sure camera is at good height and angle. Mark the beam position on camera by drawing a circle for the laser beam shown on TV screen.
3. Change back to few milliwatt. Putting OPO crystal part on rotation stage with HR side facing laser. Checking injection/reflection overlapping and transmission reaching the circle on TV screen. If not, adjusting by hand the position of OPO crystal part.
4. Adjusting camera lens to zoom in and inspect the image of crystal transmission. This image should be a circular spot without any clipping. If not, consider to adjust the position of OPO crystal part in transversal, longitudinal and tilt DOF.
5. Fixing OPO crystal part on rotation stage gently. In my case, the OPO didn't move after this. If it is moved a bit, bring it back by adjusting rotation stage. Be careful that the longer side of rotation stage should be aligned with laser.
6. Putting in-coupling mirror part of OPO and half-fixing it. Half-fixing means that the screws are ~20 degrees before the feeling of tight screws. This is to make sure that we can adjust in-coupling mirror while keeping it not going to far after each adjustment.
7. Zooming out camera to find roughly the flash of OPO cavity. When zooming out, we can see the inner cavity reflection more clearly.
8. Moving incoupling mirror by hand in horizontal direction to make inner cavity reflection good for horizontal. Moving vertical adjusting screws to find flash.
9. Zooming in camera to see the Hermite-Gaussian mode of OPO cavity. The scan speed was set to be 3Hz.
10. Adjuting finely the incoupling mirror position until the mode shape looks to be good enough. In the end, maybe there is still some residual modes. But it will be fine.
11. Change power to few hundred of milliwatt. Putting a BS before camera and direct the light to a PD to see cavity scan spectrum on an oscilloscope.
12. Fixing in-coupling mirror by the sequence of diagonal so that the alignment can be less affected.
13. Adjusting the injection laser to finalize the alignment.
After doing this procedure, we got OPO cavity scan spectrum as Fig.1. The zoom-in of TEM00 is Fig.2. We haven't characterized how much power is on higher order modes. But it looks decent for us.
The camera used in this set-up is found to be tilted today. Fig.1 shows that this tilt is introduced through the interface between camera and post.
During the alignment, we use camera to characterize cavity HG modes shapes. When the beam and camera height is aligned, we found beam appears not in the center of screen as Fig.2. We suspect the tilt between camera and post is the cause of the cavity modes mis-center as Fig.2.
Note that Fig.2 shows the non-ideal alignment of crystal, which has a black defect.

Yuhang and Michael
We checked the incoupling mirror again after Yuhang applied and peeled First Contact on each day of the holidays. It seems to be much cleaner now. Perhaps the light spots that can still be seen are internal scattering or something like that.

Here is the result of the vertical and horizontal angles of incidence that will be used during the measurement :
vertical ; 4.75 mrad or 0.272 deg
horizontal ; 41.94 mrad or 2.403 deg

Here is the result of spare ETMY bulk absorption with the 2 ears flats.
The measured value agrees with the one measured at Caltech and reported in entry 1583

Abe, Marc and help from Yuhang to move safely the mirror
We removed the spare ETMY but did not placed it inside its box as we might put first contact before the shipping back to ICRR.
We checked the surface and bulk calibration without tuning the pump nor the probe beams alignment and got R_surface = 17.87 /W and R_bulk = 0.6503 cm/W
We measured the pump beam incidence angle and installed the SHINKOSHA 7 on the translation stage (X_center = 398 mm, Y_center = 122 mm and Z_center = 71 mm).
After the usual calibration we started the measurement with s polarization at the input.
Here is the result of the vertical and horizontal angles of incidence that will be used during the measurement :
vertical ; 4.75 mrad or 0.272 deg
horizontal ; 41.94 mrad or 2.403 deg